• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Hall A Techs

    Log entry time 16:22:36 on March 24, 2014

    Entry number 387464


    Beam line valves operational

    Beam line valves from Compton to the dump are operational

    The target, spectrometer entrance, spectrometer exit and VDC covers are removed Angle limits for the left HRS is 15 degrees and locked.

    Angle limits for the right HRS is 15 degrees and locked.

    The left and right snout collimators have been removed. Collimators are in the open position.

    There appears to be a solid target in the beam position

    Both spectrometer magnets are operational with the exception of the right Q1 magnet

    Tech on call is Heidi Fansler.